# Towards White Box Deep Learning

### v4

@@ Abstract
This paper introduces semantic features as a candidate conceptual framework for white box neural networks. The proof of concept model is well-motivated, inherently interpretable, has low parameter-count and achieves almost human-level adversarial test metrics - with no adversarial training! These results and the general nature of the approach warrant further research on semantic features. The code is available at https://github.com/314-Foundation/white-box-nn.
@@ 1 Introduction
The core advantages of deep neural networks (DNNs) are their architectural simplicity and automatic feature learning. The latter is crucial for working with unstructured data as developers don’t need to design features by hand. However giving away the control over features leads to black box models - DNNs tend to learn uninterpretable "shortcut" correlations [13] that leak from train to test [15], hampering alignment and out-of-distribution performance. In particular, this gives rise to adversarial attacks [30] - semantically negligible perturbations of data that arbitrarily change model’s predictions. Adversarial vulnerability is a widespread phenomenon (vision, segmentation, NLP, sound, RL, tabular data) and largely contributes to the general lack of trust in DNNs, substantially limiting their adoption in high-stakes applications such as healthcare, military, autonomous vehicles or cybersecurity. Conversely, the main advantage of hand-designed features is the fine-grained control over model’s performance; however, such systems quickly become infeasibly complex.
@@
Figure 1: Visualisation of semantic feature $f_{\mathbf{P}\mathbf{L}}$ and its matching mechanism (SFmatch, see Section 4.2). The semantic feature $f_{\mathbf{P}\mathbf{L}}$ consists of a base feature $f$ and a set of its "small" modifications $\mathbf{L}(p_{i})(f)$. In this case, the base feature has the same dimensions as the input image and its modifications are 2D affine transformations. $\mathrm{SFmatch}(d,f_{\mathbf{P}\mathbf{L}})$ takes the maximum of the dot products $<d,\mathbf{L}(p_{i})(f)>$ thus identifying all $\mathbf{L}(p_{i})(f)$ with $f$. Affine transformations are typically parameterized by extended matrices of dimension $2\times 3$; both the base feature and the parameters $p_{i}$ of its modifications are learned. If $f$ and its modifications are easily understood by humans, then the neural network layer composed of such features can be considered a white box layer. The precise definition of semantic feature is found in Section 4.
@@ 2.1 Adversarial robustness
To address this issue the field of certified robustness has emerged with the goal of providing theoretically certified upper bounds for adversarial error [19]. However the complete certification faces serious obstacles as it is NP-Complete [33] and therefore limited to small datasets and networks, and/or to probabilistic approaches. Another shortcoming is that such guarantees can only be given for a precisely defined threat model which is usually assumed to be a $L_{p}$-bounded perturbation. The methods, results and theoretical obstacles can be vastly different222for example SOTA certified robustness for MNIST is ~94% for $L_{\infty},eps=0.3$ and only ~73% for $L_{2},eps=1.58$, see https://sokcertifiedrobustness.github.io/leaderboard/. for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards chosen $L_{p}$-certification.
@@ 2.2 Explainability
Recent work indicates a link between adversarial robustness and inherent interpretability as adversarially trained models tend to enjoy Perceptually-Aligned Gradients (PAG) - taking gradient steps along the input to maximise the activation of a selected neuron in the representation space often yields visually aligned features [10, 17]. There is also some evidence for the opposite implication that PAG implies certain form of adversarial robustness [12, 29]. However Section 6.4 of this paper shows that the latter implication does not hold - model can be inherently interpretable but not robust.
Susceptibility to adversarial attacks is arguably the main drawback of known approaches to Explainable AI (XAI). One would reasonably expect explainable models to rely exclusively on human-aligned features instead of spurious "shortcut" signals [13] and therefore be robust to adversaries. However the post-hoc attribution methods, which aim to explain the decisions of already-trained models, are fragile [1] and themselves prone to attacks [14]. On the other hand the inherently interpretable solutions such as Prototypical Parts [20, 8], Bag-of-Features [5], Self-Explaining Networks [2], SITE [32], Recurrent Attention [11], Neurosymbolic AI [28] or Capsule Networks [25] all employ various classic black box encoder networks and thus introduce the adversarial vulnerability into their cores.
@@ 3 Methodology
This paper focuses on the simplest relevant problem (Minimum Viable Dataset, MVD) to cut off the obfuscating details and get into the core of what makes neural networks fragile and uninterpretable. Despite much effort, DNNs remain susceptible to adversarial attacks even on MNIST [27] and therefore cannot be considered human-aligned, see Figure 2. That being said, the MVD for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [31] and might be considered harder than the entire MNIST in terms of average adversarial accuracy. For completeness, the preliminary results on the full MNIST are discussed in Appendix C.
@@ 4 Semantic features
  • $\mathbf{L}$ locality function : $\mathbf{P}\rightarrow\textrm{Aut}(\mathbb{R}^{s},\mathbb{R}^{s})$ where Aut is the family of differentiable automorphisms of $\mathbb{R}^{s}$
@@
Intuitively, semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [25]. Thus $(f,\mathbf{P},\mathbf{L})$ is locality-sensitive in the adequate topology capturing semantics of the domain. The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively. The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias333by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias. for the given modality.
@@ 4.1 Examples of semantic features
    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel444can be made differentiable as in [23, 16].
@@
    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices555preferably close to the identity matrix.
    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images4
@@ 4.2 Matching semantic features
The role of semantic features is to be matched against the input to uncover the general structure of the data. Suppose that we have already defined a function match: $(\mathbb{R}^{c},\mathbb{R}^{s})\rightarrow\mathbb{R}$ that measures an extent to which datapoint $d\in\mathbf{R}^{c}$ contains a feature $g\in\mathbb{R}^{s}$. Now let’s define
@@
If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$ and in this sense semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$. On the other hand every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates666i.e. $f_{p}$ is a conjunction of lower-level features.. Thus semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be used by neural networks. This is particularly apparent in the context of logical semantic features which can express logical sentences explicitly, e.g. sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
@@
  • for convolutional777an alternative definition would treat every $g^{i}$ as a separate semantic feature; we stick to the max option for simplicity. $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$
  • for affine $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot\frac{g}{||g||_{2}}$ where $g$ is a 2D image of shape $(c_{in},k,k)$ and $d$ is a 2D image patch of the same shape
@@ 5.1 Two Step Layer
MNIST datapoint is a $28\times 28$ grayscale image. Its basic building block is a pixel $x\in[0,1]$. Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH888the MEH state is an undecided state in-between.. Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
@@ 5.2 Convolutional Semantic Layer
The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$. This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with the $5\times 5$ neighbourhood of $(x,y)$ and takes the maximum across all the matches999a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$.. Intuitively, this layer checks if pixels have adequately structured neighbourhoods.
@@ 5.3 Affine Layer
The basic semantic building blocks of MNIST digits are fragments of various lines together with their rough locations in the image (same line at the top and the bottom often has different semantics). In short they are shapes at locations. The semantic identity of shape is not affected by small101010close to the identity augmented matrix. affine transformations. The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
@@ 6.1 Training
The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$. The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability111111mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it.. Those parameters were chosen "by hand" after few rounds of experimentation. The results seem pretty robust to the selection of hyperparameters as long as they are within reasonable bounds. Training a single epoch on a single CPU takes around 9s.
@@ 7.2 Limitations
Semantic features require locality engineering as they rely on the adequate domain-specific definition of the locality function. Therefore they won’t be applicable in contexts where such locality is hard to define. On the other hand it opens up interesting areas for research.
The Logical Layer has been developed as an attempt to formalize the idea of how the generic hidden layer activation pattern should look like; however it might turn out to be prone to the "no free lunch" theorem (e.g. on MNIST the properly-initialized linear probe seems to be sufficient).
@@ 8 Further research
Future work would have to involve more complex datasets and more nuanced semantic features. For example notice the analogy between affine semantic features and the SIFT [21] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while previous layers (Two Step and Convolutional Semantic) are essentially performing input normalization. This might be a helpful analogy for designing more advanced semantic features that sample across varying scales, sharpness or illumination levels.
There might be a concern that adding more layers to the network would reduce its interpretability. However the role of additional layers would be to widen the receptive field of the network. The Affine Layer defined earlier matches shapes-at-locations of center-aligned objects. Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the Global Affine Layer matches objects-at-locations which form a scene (a higher-level object). This would in fact be similar to applying Affine Layer to the "zoomed-out" version of the input.
The ablations in Section 6.4 suggest that adversarial vulnerability of DNNs might be mostly due to the fragility of the initial layers. Therefore another direction forward would be to re-examine the known approaches to XAI and try to "bug fix" their encoder networks in similar fashion.
@@
  • study non-affine perturbations of features, e.g. projective perturbations

### v5

@@ Abstract
Deep neural networks learn fragile "shortcut" features, rendering them difficult to interpret (black box) and vulnerable to adversarial attacks. This paper proposes semantic features as a general architectural solution to this problem. The main idea is to make features locality-sensitive in the adequate semantic topology of the domain, thus introducing a strong regularization. The proof of concept network is lightweight, inherently interpretable and achieves almost human-level adversarial test metrics - with no adversarial training! These results and the general nature of the approach warrant further research on semantic features. The code is available at https://github.com/314-Foundation/white-box-nn.
@@ 1 Introduction
The main advantages of deep neural networks (DNNs) are their architectural simplicity and automatic feature learning. The latter is crucial for working with unstructured data as developers don’t need to design features by hand. However, giving away the control over features leads to black box models - DNNs tend to learn hardly interpretable "shortcut" correlations [17] that leak from train to test [20], hampering alignment and out-of-distribution performance. In particular, this gives rise to adversarial attacks [35] - semantically negligible perturbations of data that arbitrarily change model’s predictions. Adversarial vulnerability is a widespread phenomenon (vision [35], segmentation/detection [39], speech recognition [9], tabular data [10], RL [19], NLP [41]) and largely contributes to the general lack of trust in DNNs, substantially limiting their adoption in high-stakes applications such as healthcare, military, autonomous vehicles or cybersecurity. Conversely, the main advantage of hand-designed features is the fine-grained control over model’s performance; however, such systems quickly become infeasibly complex.
@@
Figure 1: Visualisation of semantic feature $f_{\mathbf{P}\mathbf{L}}$ and its matching mechanism (SFmatch, see Section 4.2). The semantic feature $f_{\mathbf{P}\mathbf{L}}$ consists of a base feature $f$ and a set of its "small" modifications $\mathbf{L}(p_{i})(f)$. In this case, the base feature has the same dimensions as the input image and its modifications are 2D affine transformations. $\mathrm{SFmatch}(d,f_{\mathbf{P}\mathbf{L}})$ takes the maximum of the scalar products $d\cdot\mathbf{L}(p_{i})(f)$ thus identifying all $\mathbf{L}(p_{i})(f)$. Affine transformations are typically parameterized by extended matrices of dimension $2\times 3$; both the base feature and the parameters $p_{i}$ of its modifications are learned. If $f$ and its modifications are easily understood by humans, then the neural network layer composed of such features can be considered a white box layer. The precise definition of semantic feature is found in Section 4.
@@ 2.1 Adversarial robustness
To address this issue the field of certified robustness has emerged with the goal of providing theoretically certified upper bounds for adversarial error [24]. However the complete certification faces serious obstacles as it is NP-Complete [38] and therefore limited to small datasets and networks, and/or to probabilistic approaches. Another shortcoming is that such guarantees can only be given for a precisely defined threat model which is usually assumed to be a $L_{p}$-bounded perturbation. The methods, results and theoretical obstacles can be vastly different222for example SOTA certified robustness for MNIST is ~94% for $L_{\infty},eps=0.3$ and only ~73% for $L_{2},eps=1.58$, see https://sokcertifiedrobustness.github.io/leaderboard/. for different $p\in\{0,1,2,\ldots,\infty\}$ and often require training the model specifically towards the chosen $L_{p}$-certification.
@@ 2.2 Explainability
Recent work indicates a link between adversarial robustness and inherent interpretability as adversarially trained models tend to enjoy Perceptually-Aligned Gradients (PAG) - taking gradient steps along the input to maximise the activation of a selected neuron in the representation space often yields visually aligned features [13, 22]. There is also some evidence for the opposite implication that PAG implies certain form of adversarial robustness [15, 34]. However, Section 6.4 of this paper shows that the latter implication does not hold - model can be inherently interpretable but not robust.
Susceptibility to adversarial attacks is arguably the main drawback of known approaches to Explainable AI (XAI). One would reasonably expect explainable models to rely exclusively on human-aligned features instead of spurious "shortcut" signals [17] and therefore be robust to adversaries. However, the post-hoc attribution methods, which aim to explain the decisions of already-trained models, are fragile [1] and themselves prone to attacks [18]. On the other hand, the inherently interpretable solutions such as Prototypical Parts [25, 11], Bag-of-Features [5], Self-Explaining Networks [2], SITE [37], Recurrent Attention [14], Neurosymbolic AI [33] or Capsule Networks [30] all employ various classic black box encoder networks and thus introduce the adversarial vulnerability into their cores.
@@
2.3 Geometric Deep Learning
The idea behind semantic features resembles the framework of Geometric Deep Learning [7] or the more recent Categorical Deep Learning [16] in that semantic features are invariant to small deformations (local symmetries). However, semantic features are more practical as they focus on the local invariance and don’t require defining a formal structure on the input space; instead, they learn to sample the semantic neighbourhood of a base feature (which is also learned). Thus, the only effort required from the developer is to implement sufficient sampling mechanism, which doesn’t need to be formal or exhaustive, e.g. small affine perturbations seem to be good enough for MNIST features (see Section 6.3.3), despite the fact that not every possible local perturbation of these features is affine. Additionally, the mentioned frameworks are high-level approaches that don’t address explicitly the core issue of adversarial vulnerability.
@@ 3 Methodology
This paper focuses on the simplest relevant problem (Minimum Viable Dataset, MVD) to cut off the obfuscating details and get into the core of what makes neural networks fragile and uninterpretable. Despite much effort, DNNs remain susceptible to adversarial attacks even on MNIST [32] and therefore cannot be considered human-aligned even for this simple dataset, see Figure 2. That being said, the MVD for this paper is chosen to be the binary subset of MNIST consisting of images of "3" and "5" - as this is one of the most often confused pairs under adversaries [36] and might be considered harder than the entire MNIST in terms of average adversarial accuracy. For completeness, the preliminary results on the full MNIST are discussed in Appendix C.
@@ 4 Semantic features
  • $\mathbf{L}$ locality function : $\mathbf{P}\rightarrow\textrm{Diff}(\mathbb{R}^{s})$ where Diff is the family333informally, $\mathbf{L}(\mathbf{P})$ should consist of ”small” diffeomorphisms. of differentiable automorphisms of $\mathbb{R}^{s}$
@@
Intuitively, semantic feature consists of a base vector $f$ and a set of it’s "small" variations $f_{p}=\mathbf{L}(p)(f)$ for all $p\in\mathbf{P}$; these variations may be called poses after [30]. Thus $(f,\mathbf{P},\mathbf{L})$ is locality-sensitive in the adequate topology capturing semantics of the domain. The differentiability of $\mathbf{L}(p)$ and of $\mathbf{L}$ itself allow for gradient updates of $f$ and $\mathbf{P}$ respectively. The intention is to learn both $f$ and $\mathbf{P}$ while defining $\mathbf{L}$ explicitly as an inductive bias444by the ”no free lunch” theorem a general theory of learning systems has to adequately account for the inductive bias. for the given modality.
@@ 4.1 Examples of semantic features
    • $\mathbf{L}$ function that maps a rotation matrix to the respective rotation of a 2D kernel555can be made differentiable as in [28, 21].
@@
    • $\mathbf{P}$ set of $2\times 3$ invertible augmented matrices666preferably close to the identity matrix.
    • $\mathbf{L}$ function that maps a $2\times 3$ matrix to the respective 2D affine transformation of images5
@@ 4.2 Matching semantic features
The role of semantic features is to be matched against the input to uncover the general structure of the data. Suppose that we have already defined a function match: $(\mathbb{R}^{c},\mathbb{R}^{s})\rightarrow\mathbb{R}$ that measures an extent to which datapoint $d\in\mathbb{R}^{c}$ contains a feature $g\in\mathbb{R}^{s}$. Now let’s define
@@
If $\mathbf{L}$ captures domain topology adequately then we may think of matching $f_{\mathbf{P}\mathbf{L}}$ as a form of local inhibition along $\mathbf{L}_{\mathbf{P}}(f)=\{L(p)(f),p\in P\}$ and in this sense semantic feature defines a XOR gate over $\mathbf{L}_{\mathbf{P}}(f)$. On the other hand, every $f_{p}\in\mathbf{L}_{\mathbf{P}}(f)$ can be viewed as a conjunction (AND gate) of its non-zero coordinates777i.e. $f_{p}$ is a conjunction of lower-level features.. Thus, semantic features turn out to be a natural way of expressing logical claims about the classical world - in the language that can be used by neural networks. This is particularly apparent in the context of logical semantic features, which can express logical sentences explicitly, e.g. sentence $(A\vee B)\wedge\neg C\wedge\neg D$ can be expressed by a logical $f_{\mathbf{P}\mathbf{L}}$ where:
@@
  • for convolutional888an alternative definition would treat every $g^{i}$ as a separate semantic feature; we stick to the max option for simplicity. $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=\max_{0<=i<c_{out}}(d*\frac{g^{i}}{||g^{i}||_{2}})$ where $g$ is a 2D convolutional kernel of shape $(c_{out},c_{in},k,k)$, $g^{i}$ is the $i$-th convolutional filter of $g$, $d$ is an image patch of shape $(c_{in},k,k)$
  • for affine $f_{\mathbf{P}\mathbf{L}}$: $\mathrm{match}(d,g)=d\cdot\frac{g}{||g||_{2}}$ where $g$ is a 2D image of shape $(c_{in},k,k)$ and $d$ is a 2D image patch of the same shape (treated as vectors in $\mathbb{R}^{c}$)
@@ 5.1 Two Step Layer
MNIST datapoint is a $28\times 28$ grayscale image. Its basic building block is a pixel $x\in[0,1]$. Despite being allowed to assume any value between 00 and $1$ it is actually (semantically) a ternary object - it’s either ON, OFF or MEH999the MEH state is an undecided state in-between.. Therefore the semantic space should squash the $[0,1]$ interval into those 3 values.
@@ 5.2 Convolutional Semantic Layer
The layer performs $\mathrm{SFmatch}(d(x,y),f_{\mathbf{P}\mathbf{L}})$ with every pixel $(x,y)$ of image $d$. This means that it matches the two rotated filters $g^{0}$ and $g^{1}$ with the $5\times 5$ neighbourhood of $(x,y)$ and takes the maximum across all the matches101010a single number, not a pair of numbers, as in the chosen definition of match for convolutional $f_{\mathbf{P}\mathbf{L}}$.. Intuitively, this layer checks if pixels have adequately structured neighbourhoods.
@@ 5.3 Affine Layer
The basic semantic building blocks of MNIST digits are fragments of various lines together with their rough locations in the image (same line at the top and the bottom often has different semantics). In short they are shapes at locations. The semantic identity of shape is not affected by small111111close to the identity augmented matrix. affine transformations. The next shape-extracting layer will therefore consist of 8 affine semantic features $(f,\mathbf{P},\mathbf{L})$ defined as follows:
@@ 6.1 Training
The model was trained to optimize the CrossEntropyLoss for 15 epochs with Adam optimizer, learning rate $3e{-3}$, weight decay $3e{-6}$. The only augmentation was Gaussian noise $\sim\mathcal{N}(0.05,0.25)$ applied with $0.7$ probability121212mean is slightly positive because of the sparsity of data - subtracting pixel value is on average more harmful than adding it.. Those parameters were chosen "by hand" after few rounds of experimentation. The results seem pretty robust to the selection of hyperparameters as long as they are within reasonable bounds. Training a single epoch on a single CPU takes around 9s.
@@ 7.2 Limitations
Semantic features require locality engineering as they rely on the adequate domain-specific definition of the locality function. Therefore, they won’t be applicable in contexts where such locality is hard to define. On the other hand, it opens up interesting areas for research.
The Logical Layer has been developed as an attempt to formalize the idea of how the generic hidden layer activation pattern should look like; however, it might turn out to be prone to the "no free lunch" theorem (e.g. on MNIST the properly-initialized linear probe seems to be sufficient).
@@ 8 Further research
Notice the analogy between affine semantic features and the SIFT [26] algorithm - the learnable pose sampling may be interpreted as finding Keypoints and their Descriptors, while the initial 2 layers of the PoC network essentially perform input normalization. This might be a helpful analogy for designing more advanced semantic features that sample across varying scales, sharpness or illumination levels.
There might be a concern that adding more layers to the network would reduce its interpretability. However, the role of additional layers would be to widen the receptive field of the network. The Affine Layer defined earlier matches shapes-at-locations of center-aligned objects. Adding additional layer would involve a sliding window approach (or Keypoint Detection similar to that in SIFT) to match center-aligned objects inside the window and integrate this with the approximate global location of the window; then the "Global" Affine Layer would match objects-at-locations forming a scene (a higher-level object). This would in fact be similar to applying Affine Layer to the "zoomed-out" input.
The ablations in Section 6.4 suggest that adversarial vulnerability of DNNs might be mostly due to the fragility of the initial layers. Therefore, another direction forward would be to re-examine the known approaches to XAI and try to "bug fix" their encoder networks in similar fashion.
@@
  • study non-affine perturbations of spatial features, e.g. projective perturbations